MICROSOLVATION OF PARTIALLY BOUND NITRILE - SO_{3} COMPLEXES

\underline{D.L. FIACCO}, M.E. OTT, S.R. WEERS, T. HOWE, Department of Chemistry, University of Minnesota, 207 Pleasant St., SE, Minneapolis, MN 55455; A. ROBERTS, Department of Chemistry, Carleton College, Northfield, MN 55057; and K.R. LEOPOLD, Department of Chemistry, University of Minnesota, 207 Pleasant St., SE, Minneapolis, MN 55455.

The rotational spectrum of HCN–HCN-SO_{3} and its deuterated isotopomers have been observed by pulsed nozzle Fourier transform microwave spectroscopy. The complex is a symmetric rotor with the atomic ordering indicated and preliminary analysis yields an S-N bond length of 2.43 Å. This distance is approximately 0.15 Å shorter than the corresponding distance in HCN-SO_{3}, indicating that the remote HCN has a large effect on the nitrogen - sulfur bond. This result represents the first direct probe of the effect of a nearest neighbor on the structure of a partially bonded adduct. We also present measurements of the dipole moments of HCN-SO_{3} and CH_{3}CN-SO_{3} (4.416 ± 0.029 D and 6.06 ± 0.18 D, respectively). Examination of a series of SO_{3} complexes indicates a substantial increase in polarity in these systems as dative bond formation proceeds. These results will be discussed in connection with the origins of the large gas-to- solid structure changes normally observed for partially bound species.